# AD-GS: Alternating Densification for Sparse-Input 3D Gaussian Splatting

### v1

@@ Abstract.
3D Gaussian Splatting (3DGS) has shown impressive results in real-time novel view synthesis. However, it often struggles under sparse-view settings, producing undesirable artifacts such as floaters, inaccurate geometry, and overfitting due to limited observations. We find that a key contributing factor is uncontrolled densification, where adding Gaussian primitives rapidly without guidance can harm geometry and cause artifacts. We propose AD-GS, a novel alternating densification framework that interleaves high and low densification phases. During high densification, the model densifies aggressively, followed by photometric loss based training to capture fine-grained scene details. Low densification then primarily involves aggressive opacity pruning of Gaussians followed by regularizing their geometry through pseudo-view consistency and edge-aware depth smoothness. This alternating approach helps reduce overfitting by carefully controlling model capacity growth while progressively refining the scene representation. Extensive experiments on challenging datasets demonstrate that AD-GS significantly improves rendering quality and geometric consistency compared to existing methods.
@@ 3.3.2. Alternating Densification Phases
Tanks & Temples
[…]
LLFF
[…]
@@ 4.3. Results
Figs. 4, 5, 6, 8, 9, 10, 11 and 12 show the qualitative results across various scenes from all datasets. AD-GS is able to recover fine-scale textures and detailed structures that other methods tend to oversmooth or miss entirely. It maintains sharper object boundaries and better geometric fidelity, particularly noticeable in scenes with complex edges and discontinuities. Moreover, AD-GS suppresses floaters and noisy artifacts more effectively, resulting in better reconstructions. Its surface reconstructions also exhibit higher smoothness and realism, especially in areas with gradual depth variations. Thus our alternating training strategy offers benefits across a diverse range of sparse-view scenarios.

### v2

@@ Abstract.
3D Gaussian Splatting (3DGS) has shown impressive results in real-time novel view synthesis. However, it often struggles under sparse-view settings, producing undesirable artifacts such as floaters, inaccurate geometry, and overfitting due to limited observations. We find that a key contributing factor is uncontrolled densification, where adding Gaussian primitives rapidly without guidance can harm geometry and cause artifacts. We propose AD-GS, a novel alternating densification framework that interleaves high and low densification phases. During high densification, the model densifies aggressively, followed by photometric loss based training to capture fine-grained scene details. Low densification then primarily involves aggressive opacity pruning of Gaussians followed by regularizing their geometry through pseudo-view consistency and edge-aware depth smoothness. This alternating approach helps reduce overfitting by carefully controlling model capacity growth while progressively refining the scene representation. Extensive experiments on challenging datasets demonstrate that AD-GS significantly improves rendering quality and geometric consistency compared to existing methods. The source code for our model can be found on our project page: https://gurutvapatle.github.io/publications/2025/ADGS.html.
@@ 3.2. Analysis of Densification under Sparse Input
Tanks & Temples
Model PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$ 3-view 6-view 9-view 3-view 6-view 9-view 3-view 6-view 9-view 3DGS 16.992 22.393 24.492 0.556 0.788 0.850 0.352 0.182 0.141 FSGS 19.644 26.436 28.488 0.637 0.847 0.888 0.312 0.168 0.139 CoR-GS 19.246 26.273 28.489 0.650 0.849 0.891 0.342 0.175 0.132 DropGaussian 19.475 26.452 28.488 0.652 0.852 0.891 0.320 0.173 0.138 AD-GS (Ours) 19.724 27.139 28.964 0.678 0.867 0.902 0.285 0.142 0.113
LLFF
Model PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$ 3-view 6-view 9-view 3-view 6-view 9-view 3-view 6-view 9-view 3DGS 18.90 22.57 24.02 0.632 0.757 0.799 0.269 0.182 0.151 FSGS 19.50 23.25 24.56 0.655 0.776 0.815 0.269 0.185 0.151 CoR-GS 19.59 23.33 24.75 0.674 0.780 0.817 0.271 0.187 0.152 DropGaussian 19.72 23.52 24.89 0.674 0.786 0.824 0.266 0.182 0.148 AD-GS (Ours) 20.06 23.54 25.02 0.699 0.793 0.830 0.237 0.170 0.142
@@ 3.3.2. Alternating Densification Phases
Mip-NeRF360
Model PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$ 12-view 24-view 12-view 24-view 12-view 24-view 3DGS 17.36 22.06 0.496 0.701 0.403 0.253 FSGS 18.51 23.11 0.547 0.720 0.411 0.275 CoR-GS 19.42 23.20 0.579 0.728 0.410 0.272 DropGaussian 19.18 23.28 0.575 0.732 0.412 0.277 AD-GS(Ours) 19.66 23.68 0.593 0.750 0.386 0.249
@@ 4.3. Results
Figs. 4, 9, 5, 7, 8, 10, 11 and 6 show the qualitative results across various scenes from all datasets. AD-GS is able to recover fine-scale textures and detailed structures that other methods tend to oversmooth or miss entirely. It maintains sharper object boundaries and better geometric fidelity, particularly noticeable in scenes with complex edges and discontinuities. Moreover, AD-GS suppresses floaters and noisy artifacts more effectively, resulting in better reconstructions. Its surface reconstructions also exhibit higher smoothness and realism, especially in areas with gradual depth variations. Thus our alternating training strategy offers benefits across a diverse range of sparse-view scenarios.